OpenStudy (anonymous):

Use the binomial theorem to find expressions in terms of r1 and r2 for: i. (x^3+y^3) ii. (x^4+y^4) r1=(x+y) and r2=(xy)

OpenStudy (anonymous):

could you help me with the working? I have no idea how to get the answer :(

OpenStudy (mokeira):

do you know \[(x ^{3}+y ^{3})=(x+y)(x ^{2}+y ^{2}-xy)\]

OpenStudy (mokeira):

do you know that formula?

OpenStudy (anonymous):

yes, but i have to use the binomial theorem, for ii

OpenStudy (mokeira):

@midhun.madhu1987

OpenStudy (mokeira):

@aryandecoolest

OpenStudy (mokeira):

@agreene

OpenStudy (mokeira):

@ganeshie8

OpenStudy (mokeira):

@hopelovelift

OpenStudy (anonymous):

1. (r1)(r1^2-3r2)

OpenStudy (anonymous):

2. r1^4-4r2( r1^2-2r2)-6r2^2

OpenStudy (anonymous):

ok that looks good thanks :) do you have any working? I would really like to know how you worked it out :)

OpenStudy (anonymous):

well yeah i have working.!!!

OpenStudy (anonymous):

hope first one is clear to you. right?

OpenStudy (anonymous):

yeah the fist one is clear, but I do not know how to use the binomial theorem, could you post the working? thanks

OpenStudy (anonymous):

\[(x+y)^4=4C _{0}*x^4+4C _{1}*x^3y+4C _{2} *x^2y^2+4C _{3} *xy^3+4C _{4}*y^4\] \[x^4+y^4=(x+y)^4-4x^3y-4xy^3-6(xy)^2\] substitute and get the answer )

OpenStudy (anonymous):

thanks! :)

OpenStudy (anonymous):

np ) anytime